# DRP: Distilled Reasoning Pruning with Skill-aware Step Decomposition for Efficient Large Reasoning Models

### v2

@@ Abstract
While Large Reasoning Models (LRMs) have demonstrated remarkable success in complex reasoning tasks through Long Chain-of-Thought (CoT) reasoning, their inference often involves excessively verbose reasoning traces, resulting in substantial inefficiency. To address this issue, we propose Distilled Reasoning Pruning (DRP), a hybrid framework that combines inference-time pruning with tuning-based distillation—two widely used strategies for efficient reasoning. DRP employs a teacher model to perform skill-aware step decomposition and content pruning, and then distills the pruned reasoning paths into a student model, enabling it to reason both efficiently and accurately. Across a series of challenging mathematical reasoning datasets, we find models trained with DRP achieve substantial improvements in token efficiency without sacrificing accuracy. Specifically, DRP reduces the average token usage on GSM8K from 917 to 328 while improving accuracy from 91.7% to 94.1%, and achieves a 43% token reduction on AIME with no performance drop. Further analysis reveals that aligning the reasoning structure of training CoTs with the student’s reasoning capacity is critical for effective knowledge transfer and performance gains. We release our code at: https://github.com/YuxuanJiang1/DRP
@@ 1 Introduction
Existing solutions predominantly follow two paradigms: inference-time pruning, which attempts to terminate generation early to avoid redundant reasoning steps Cui et al. (2025); Muennighoff et al. (2025); Fu et al. (2024); Zeng et al. (2025), and distillation-based compression, where smaller models are trained on teacher-generated reasoning paths to imitate the concise reasoning behavior of larger models Tan et al. (2024); Zhu et al. (2024); Xu et al. (2025b). However, both approaches can be at the cost of accuracy: pruning methods risk prematurely halting the reasoning process, thereby preventing the model from reaching the correct answer, while distillation methods tend to underperform—particularly for small-scale student models—due to the learnability gap, which hinders effective knowledge transfer when the performance or size gap between student and teacher is too great Xu et al. (2024); Li et al. (2025b).
@@
Figure 1: An overview of our proposed Distilled Reasoning Pruning framework, which unifies pruning and distillation by having the teacher model prune the student model’s reasoning chains and distill the resulting concise CoT back into the student. This approach improves reasoning efficiency without sacrificing accuracy.
Figure 2: Overview of our proposed DRP framework, which distills concise, skill-aware reasoning paths into a target model to improve reasoning efficiency while preserving accuracy.
@@
To address these limitations, we propose Distilled Reasoning Pruning (DRP) (Fig. 1), a hybrid framework that combines the strengths of both pruning and distillation (student/teacher) paradigms. In particular, rather than simply distilling down a response from the teacher model, we use the teacher model to prune an initial, lengthy, CoT reasoning trajectory from the student model. To facilitate this, and to encourage shorter yet informative resulting pruned trajectories, we introduce a skill-based step decomposition method. The teacher model uses this to prune the trajectory, which produces more stable and semantically coherent reasoning units. Unlike prior methods that rely solely on either teacher-generated Xu et al. (2025b); Zhu et al. (2024) or self-sampled Chen et al. (2024); Ma et al. (2025) concise trajectories for distillation, DRP takes advantage of the teacher’s pruning within the student model’s original reasoning structure. This design reduces the learnability gap and enables student models to achieve efficient reasoning without compromising performance.
@@
Consider the example in Fig. 2. For this word problem, the student model generates its initial CoT trajectory (shown in the <think> block). This trajectory is provided to a teacher model, which segments that CoT into steps, with a high-level description of the skill that that step demonstrates or exercises. The teacher model then prunes and curates these skill-segmented steps, such as by merging similar ones (if the student was redundant or verbose) or deleting steps (e.g., for backtracking). This pruned CoT trajectory is then provided to the student model for supervised fine-tuning.
@@ 2.1 The Overthinking Problem
Large reasoning models, such as OpenAI o1 OpenAI (2024b) and DeepSeek-R1 Guo et al. (2025), are a specialized subclass of large language models trained to iteratively generate intermediate reasoning steps and refine the thinking process until reaching a final answer Chen et al. (2025); Sui et al. (2025); Yu et al. (2025). Unlike traditional LLMs, which rely on test-time strategies such as Chain-of-Thought prompting to elicit detailed reasoning Tong et al. (2024), LRMs are trained to internalize CoT generation. This design significantly enhances their performance on reasoning-intensive tasks, particularly in challenging mathematics and programming. However, LRMs, especially those with smaller parameter sizes, often exhibit a tendency to produce overly verbose reasoning chains, including unnecessary tracebacks and redundant alternative paths Chen et al. (2024); Sui et al. (2025); Cui et al. (2025); Fu et al. (2025, 2024). Such overthinking not only increases token consumption beyond practical inference budgets but also introduces reasoning drift that can degrade output correctness Hou et al. (2025); Sui et al. (2025). To address these, our method explicitly prunes harmful overthinking behaviors while simultaneously guiding the model toward concise and effective reasoning expressions.
@@ 2.2 Token-Efficient Reasoning Methods
Current token-efficient reasoning methods can be categorized into three approaches Sui et al. (2025):
(1) Prompt-based methodsimpose token budget constraints within the prompt to encourage concise reasoning without altering model parameters. For example, TALE Han et al. (2024) estimates per-instance token budgets to reduce length while maintaining accuracy on simple tasks. However, these methods rely on handcrafted prompt designs and struggle with complex inputs. (2) Supervised fine-tuning (SFT)approaches train models on compressed reasoning traces to promote internalized efficiency Xia et al. (2025); Munkhbat et al. (2025). CoT-Valve Ma et al. (2025), for instance, uses iterative compression stages to fine-tune reasoning conciseness. Yet, these approaches often require task-specific data and retraining. (3) Reinforcement learning (RL)methods guide models toward concise reasoning by introducing explicit reward signals that penalize overly long outputs Team et al. (2025); Chen et al. (2024). Some approaches incorporate early-exit mechanisms at inference time Muennighoff et al. (2025); Fu et al. (2024); Zeng et al. (2025). A representative method is ThinkPrune Hou et al. (2025), which defines a target reasoning length and progressively fine-tune models under increasingly tighter constraints.
Despite progress, existing approaches often trade off accuracy, especially in out-of-domain (OOD) tasks, for brevity. In contrast, we propose a lightweight framework that leverages external teacher models to perform skill-aware pruning over reasoning steps. This structure-aware intervention enables significant token reduction while improving performance, especially under distribution shift.
@@ 3 Methodology
We propose Distilled Reasoning Pruning (DRP), a method that distills concise reasoning paths into a student model by first decomposing them into skill-based steps and then pruning redundant content with the help of a teacher model (e.g., GPT-4o). As illustrated in Figure 2, DRP proceeds in three stages: (1) decomposing the original reasoning into fine-grained, skill-based steps; (2) pruning and rewriting these steps via a teacher LLM; and (3) fine-tuning the student model on the revised, compact trajectories. These stages result in token-efficient supervision that improves both reasoning efficiency and accuracy during fine-tuning.
@@ 3.1 Skill-Based Step Decomposition
To solve a math problem, a reasoning model generates a token sequence as the full response, which consists of two components: a structured thinking process $T$, enclosed within the <think> </think> tags, and a final answer summarization $A$, captures the conclusive statement derived from the reasoning trace. Together, we terms these a “response,” $R=(T,A)$. For instance, in example shown in Fig. 2, the “thinking” portion of the initial response would include “Natalia sold 48 clips in …Total clips sold: 48 + 24 = 72” while the answer summarization is “The answer is 72.”
We extract the thinking segment $T$ from the student model’s response and prompt a teacher model to decompose it into a sequence of non-overlapping segments, each aligned with a functional reasoning skill such as arithmetic computation, comparison, or logical inference. This skill-based decomposition serves as the foundation for step-level pruning and distillation. Compared to naïve sentence-based splitting, it enhances the stability of step boundaries, maintains semantic coherence across reasoning steps, and provides a more fine-grained supervision signal (in § 6.1). We validate the quality of this segmentation through a pairwise evaluation using Gemini 2.0 Flash as an external judge; results and examples are presented in Appendix D for review.
Formally, given a response $R=(T,A)$, we define a decomposition function $D$ that maps the thinking trace into skill-annotated steps:
@@
where each $s_{i}$ denotes a contiguous span of tokens representing a single reasoning step, and $k_{i}$ is the associated skill label, inferred by the teacher model based on its internal knowledge and understanding of the reasoning process. For the example in Fig. 2, one segment should be “Natalia sold 48 clips in April,” which corresponding to the skill “Reading given quantity.” However, these skills can cover broader topics such as “Algebraic representation” and “Interpret a fraction of a subset.” The complete decomposition prompts are provided in Appendix C.
@@ 3.2 Step-Level Pruning
Finally, the teacher model rewrites $\hat{T}$ into a fluent, coherent reasoning trace that preserves the tone and speaking style of the student model. To ensure consistency between the revised reasoning and the final answer, we prompt the teacher to optionally revise the original answer segment $A$, yielding an updated final answer summarization $\hat{A}$. The final output becomes a concatenation $\hat{R}=(\hat{T},\hat{A})$, which we use as the target for supervised fine-tuning. Complete prompt are in Appendix C.
@@ 3.3 Supervised Fine-Tuning
We construct the training dataset using pairs $(x,\hat{R})$, where $x$ is the input question and $\hat{R}$ is the complete revised response.
@@ 4.2 Models
We use DeepSeek-R1-Distill-Qwen-7B and DeepSeek-R1-Distill-Qwen-1.5B Guo et al. (2025) as student models for supervised fine-tuning. Both are distilled variants of DeepSeek-R1 optimized for efficient inference. Our primary teacher model is GPT-4o OpenAI (2024a), which performs step-level decomposition and pruning to generate supervision signals. For ablation studies, we additionally explore alternative teacher models, including Gemini 2.0 Flash DeepMind (2025) and ChatGPT OpenAI (2022).
@@ 4.6 Evaluation Metrics
We report Pass@1 as the accuracy metric. For efficiency, we measure the number of reasoning tokens generated per completion using the HuggingFace-compatible Qwen tokenizer. [^3] https://huggingface.co/Qwen/Qwen-tokenizer
@@ 6 Ablation Studies
Dataset Method Accuracy Tokens
GSM8K 7B Base 91.7% 917
Default 92.7% 350
DRP 94.1% 328
OOD Tasks
MATH500 7B Base 92.4% 2486
Default 92.0% 1905
DRP 93.0% 1781
AIME24 7B Base 15/30 8674
Default 14/30 4678
DRP 15/30 4966
AMC 7B Base 31/40 4845
Default 31/40 4975
DRP 33/40 3258
Table 2: Comparison of 7B base model, default step segmentation, and our DRP method across in-distribution and out-of-distribution math benchmarks.
To assess the impact of our skill-based decomposition on downstream performance, we compare to a baseline that applies a default step-wise split without leveraging any functional skill information. Both variants are followed by teacher-guided revision and used to train models within the same DRP framework.
@@
As shown in Table 2, while default segmentation already provides some benefit under the self-revision setup, our skill-based segmentation consistently achieves stronger results across all benchmarks. DRP with skill-aware segmentation not only improves accuracy (e.g., +2 points on AMC) but also reduces token usage more effectively (e.g., 3258 vs. 4975 tokens on AMC).
This performance gap can be attributed in part to the difference in decomposition granularity: on the GSM8K training set, our method produces an average of 12.6 steps per example, compared to only 8.3 steps under the default strategy (see Appendix E for an illustrative example). The finer-grained, skill-aligned structure facilitates more precise pruning and better alignment with distinct math operations.
Crucially, this structural advantage also translates into improved generalization: we observe that the benefit of skill-based segmentation holds even on out-of-distribution reasoning tasks, further demonstrating its role in producing stable and transferable reasoning supervision.
@@
Table 3: Impact of different teacher models on DRP performance. We compare GPT-4o, Gemini 2.0 Flash, and ChatGPT as pruning teachers, evaluating downstream accuracy and average token usage on four math benchmarks.
@@
To answer this, we compare DRP with a direct distillation baseline, where the student model is trained on concise CoT generated directly by GPT-4o. These responses are notably short—averaging 186 tokens—compared to DRP’s CoT average of 330 tokens. Table 4 presents the results.
We observe that:
  • On the in-distribution dataset GSM8K, direct distillation reduces token usage by over 50% (425 vs. 917 tokens) with only a minor accuracy drop (90.7% vs. 91.7%), indicating its effectiveness in simpler tasks.
  • On OOD tasks like MATH500, AIME24, and AMC, however, accuracy drops significantly despite only modest token savings (e.g., 88.6% vs. 93.0% on MATH500; 13/30 vs. 15/30 on AIME24), suggesting limited generalization.
These findings suggest that shorter CoT alone do not guarantee better generalization in reasoning tasks. While direct distillation from teacher LLMs provides compressed outputs, these often omit essential intermediate reasoning steps. In contrast, DRP explicitly removes redundant content while preserving a step-level, skill-aligned structure, allowing the model to internalize more transferable problem-solving patterns.
This highlights an important insight: Retaining structurally meaningful reasoning—even at the cost of moderate length—yields superior downstream performance compared to blindly imitating shorter outputs. A comparison between short-form and DRP-style trajectories is in Appendix F.

### v3

@@ Abstract
While Large Reasoning Models (LRMs) excel at complex tasks via long Chain-of-Thought (CoT) reasoning, their outputs are often excessively verbose, leading to inefficiency. This problem is amplified when the student’s long-form reasoning mismatches the concise outputs of smaller teacher models—common in LLM distillation to avoid using costly large teachers. To address this issue, we propose Distilled Reasoning Pruning (DRP), a hybrid framework that combines inference-time pruning with tuning-based distillation. DRP leverages a teacher model to perform skill-aware step decomposition and pruning, then distills the refined reasoning paths into a student model, enabling efficient and accurate reasoning. Across challenging math datasets, DRP significantly reduces token usage without sacrificing accuracy—for instance, cutting tokens on GSM8K from 917 to 328 while improving accuracy from 91.7% to 94.1%, and reducing AIME tokens by 43% with no performance drop. Further analysis shows that aligning training CoT structure with the student’s capacity is key to effective knowledge transfer. Code is available at: https://github.com/YuxuanJiang1/DRP
@@ 1 Introduction
Figure 1: An overview of our proposed Distilled Reasoning Pruning (DRP) framework, which unifies pruning and distillation. Unlike traditional distillation, DRP uses a teacher LLM to prune the student model’s Long-CoT reasoning chains into concise CoTs, which are then distilled back into the student. This design addresses the reasoning style mismatch between verbose student models and concise teacher models, improving efficiency without sacrificing accuracy.
Existing solutions predominantly follow two paradigms: inference-time pruning, which attempts to terminate generation early to avoid redundant reasoning steps Cui et al. (2025); Muennighoff et al. (2025); Fu et al. (2024); Zeng et al. (2025), and distillation-based compression, where smaller models are trained on teacher-generated reasoning paths to imitate the concise reasoning behavior of larger models Tan et al. (2024); Zhu et al. (2024); Xu et al. (2025b). However, both approaches can be at the cost of accuracy: pruning methods risk prematurely halting the reasoning process, while distillation methods tend to underperform due to the learnability gap Xu et al. (2024); Li et al. (2025b). This gap becomes especially pronounced when the teacher adopts a Short-CoT style (concise, polished reasoning), and the student follows a Long-CoT style (verbose reasoning with reflective self-corrections). Such style mismatch introduces a compatibility issue that hinders effective learning, as discussed in prior work Xu et al. (2024).
Figure 2: Overview of DRP framework. The student model generates Long-CoT reasoning traces, which are first decomposed into skill-based steps, then pruned and refined with help from a teacher LLM (e.g., GPT-4o). The concise CoTs align better with the student’s learning, improving efficiency without sacrificing accuracy.
@@
To address these limitations, we propose Distilled Reasoning Pruning (DRP) (Fig.˜1), a hybrid framework that combines the strengths of both pruning and distillation (student/teacher) paradigms. In particular, rather than simply distilling down a response from the teacher model, we use the teacher model to prune an initial, lengthy, CoT reasoning trajectory from the student model. To facilitate this, and to encourage shorter yet informative resulting pruned trajectories, we introduce a skill-based step decomposition method. The teacher model uses this to prune the trajectory, which produces more stable and semantically coherent reasoning units. Unlike prior methods that rely solely on either teacher-generated Xu et al. (2025b); Zhu et al. (2024) or self-sampled Chen et al. (2024); Ma et al. (2025) concise trajectories for distillation, DRP takes advantage of the teacher’s pruning within the student model’s original reasoning structure. This design reduces the learnability gap and enables student models to achieve efficient reasoning without compromising performance.
@@
Consider the example in Fig.˜2. For this word problem, the student model generates its initial CoT trajectory (shown in the <think> block). This trajectory is provided to a teacher model, which segments that CoT into steps, with a high-level description of the skill that that step demonstrates or exercises. The teacher model then prunes and curates these skill-segmented steps, such as by merging similar ones (if the student was redundant or verbose) or deleting steps (e.g., for backtracking). This pruned CoT trajectory is then provided to the student model for supervised fine-tuning.
@@ 2.1 The Overthinking Problem
Large reasoning models , such as OpenAI o1 OpenAI (2024b) and DeepSeek-R1 Guo et al. (2025), are a subclass of LLMs trained to iteratively generate and refine intermediate steps Chen et al. (2025); Sui et al. (2025); Yu et al. (2025), internalizing Chain-of-Thought (CoT) reasoning Tong et al. (2024). This leads to strong performance on complex tasks in math and code. However, LRMs—especially smaller ones—often overgenerate verbose reasoning chains with unnecessary tracebacks and redundant paths Chen et al. (2024); Cui et al. (2025); Fu et al. (2025), increasing token usage and causing reasoning drift that may harm accuracy Hou et al. (2025); Sui et al. (2025).
Our method addresses this overthinking by pruning redundant steps and promoting concise, effective reasoning.
@@ 2.2 Token-Efficient Reasoning Methods
Current token-efficient reasoning methods fall into three categories Sui et al. (2025):
(1) Prompt-based methodsconstrain token budgets at the prompt level to encourage brevity without retraining. TALE Han et al. (2024), for example, estimates per-instance budgets to reduce output length. However, these methods depend on handcrafted prompts and struggle with complex tasks. (2) Supervised fine-tuning (SFT)trains models on compressed traces to internalize efficiency Xia et al. (2025); Munkhbat et al. (2025), as in CoT-Valve Ma et al. (2025). These methods typically require task-specific data and retraining. (3) Reinforcement learning (RL)introduces rewards to penalize long outputs Team et al. (2025); Chen et al. (2024), sometimes with early-exit mechanisms Muennighoff et al. (2025); Fu et al. (2024); Zeng et al. (2025); Dai et al. (2025); Yang et al. (2025). ThinkPrune Hou et al. (2025) sets target lengths and tightens constraints over time.
Despite these advances, many approaches reduce accuracy—especially on out-of-domain (OOD) tasks—when optimizing for brevity. By contrast, our framework uses external teacher models to perform skill-aware pruning, reducing tokens while improving robustness under distribution shift.
2.3 LLM Self-Refinement
Recent work explores self-refinement, where a model iteratively revises its own outputs to improve accuracy Madaan et al. (2023); Li et al. (2024b), often through self-feedback or selective re-generation.
In contrast, our approach introduces external teacher supervision to prune redundant steps during reasoning. Rather than focusing on iterative correction of final answers, our skill-aware pruning directly optimizes the reasoning structure, thereby improving both token efficiency and reasoning quality.
@@ 3 Methodology
We propose Distilled Reasoning Pruning (DRP), a method to improve the efficiency of Long-CoT student models by refining their reasoning traces using a concise Short-CoT teacher. This asymmetric setup is key: the student (e.g., R1-Distill-Qwen-7B) generates verbose, reflective reasoning, while the teacher (e.g., GPT-4o) produces polished, compact CoTs. To bridge this gap, we let the student generate initial traces, which are selectively revised under teacher guidance.
As shown in Figure 2, DRP involves three stages: (1) decomposing reasoning into fine-grained, skill-based steps; (2) pruning and rewriting via a teacher LLM; and (3) fine-tuning the student on the revised traces. This process yields token-efficient supervision that enhances both accuracy and reasoning efficiency.
@@ 3.1 Skill-Based Step Decomposition
To solve a math problem, a reasoning model produces a response consisting of two parts: a structured reasoning trace $T$ enclosed in <think> </think> tags, and a final answer summarization $A$. We denote the response as $R=(T,A)$. For example, in Fig.˜2, the trace may include “Natalia sold 48 clips in …Total clips sold: 48 + 24 = 72,” while the answer summarization is “The answer is 72.”
We extract the trace $T$ and prompt a teacher model to decompose it into non-overlapping segments, each aligned with a functional reasoning skill (e.g., arithmetic, comparison, logical inference). This skill-based segmentation supports step-level pruning and distillation. Compared with naïve sentence splitting, it yields more stable step boundaries, preserves semantic coherence, and provides finer-grained supervision signals (see §˜6.1). The segmentation quality is validated through pairwise evaluation with Gemini 2.0 Flash; results and examples appear in Appendix D.
Formally, given a response $R=(T,A)$, we define a decomposition function $D$:
@@
where each $s_{i}$ is a contiguous token span representing one reasoning step, and $k_{i}$ is the corresponding skill label assigned by the teacher model. For instance, “Natalia sold 48 clips in April” corresponds to the skill “Reading given quantity.” Skills also cover broader tasks such as “Algebraic representation” or “Interpreting fractions of a subset.” Complete decomposition prompts are provided in Appendix C.
@@ 3.2 Step-Level Pruning
[…]
@@ 3.3 Supervised Fine-Tuning
[…]
@@ 4.2 Models
We use DeepSeek-R1-Distill-Qwen-7B and DeepSeek-R1-Distill-Qwen-1.5B Guo et al. (2025) as student models for supervised fine-tuning. Both are distilled variants of DeepSeek-R1 optimized for efficient inference. Our primary teacher model is GPT-4o OpenAI (2024a), which performs step-level decomposition and pruning to generate supervision signals. For ablation studies, we additionally explore alternative teacher models, including Gemini 2.0 Flash DeepMind (2025), Deepseek V3 Liu et al. (2024) and ChatGPT OpenAI (2022).
@@ 4.6 Evaluation Metrics
We report Pass@1 as the accuracy metric, averaged over five independent runs to account for randomness in decoding. For efficiency, we measure the number of reasoning tokens generated per completion using the HuggingFace-compatible Qwen tokenizer. [^3] https://huggingface.co/Qwen/Qwen-tokenizer
@@ 5.3 Token Usage Distribution Shift
Dataset Method Accuracy Tokens
GSM8K 7B Base 91.7% 917
No decomposing 91.0% 434
Default 92.7% 350
DRP 94.1% 328
OOD Tasks
MATH500 7B Base 92.4% 2486
No decomposing 88.6% 2102
Default 92.0% 1905
DRP 93.0% 1781
AIME24 7B Base 15/30 8674
No decomposing 13/30 6201
Default 14/30 4678
DRP 15/30 4966
AMC 7B Base 31/40 4845
No decomposing 29/40 4028
Default 31/40 4975
DRP 33/40 3258
Table 2: Comparison of 7B base model, no decomposition, default step segmentation, and our DRP method across in-distribution and out-of-distribution math benchmarks.
@@ 6 Ablation Studies
To evaluate the impact of our skill-based decomposition, we compare three variants under the same DRP framework: (1) skill-based segmentation, (2) default step-wise splitting without skill labels, and (3) no decomposition at all.
@@
As shown in Table 2, skill-based segmentation consistently outperforms the other variants, improving accuracy (e.g., +2 on AMC) and reducing token usage more effectively (e.g., 3258 vs. 4975). This gain is partly due to finer granularity—our method produces 12.6 steps per example on GSM8K, compared to 8.3 with default splitting (see Appendix E).
@@
   with DeepSeek-V3 teacher 92.7% 340 90.4% 1908 14/30 4640 32/40 3136
Table 3: Impact of different teacher models on DRP performance. We compare GPT-4o, Gemini 2.0 Flash, ChatGPT, and DeepSeek-V3 as pruning teachers, evaluating downstream accuracy and average token usage.
In contrast, skipping decomposition entirely leads to significant performance degradation across all benchmarks, confirming that explicit step segmentation is essential for effective pruning. Notably, this structural benefit generalizes: skill-based segmentation remains robust even on out-of-distribution tasks, yielding more transferable supervision.
@@
To answer this, we compare DRP with a direct distillation baseline, where the student model is trained on concise CoT generated directly by GPT-4o. These responses are notably short—averaging 186 tokens—compared to DRP’s CoT average of 330 tokens. Table 4 presents the results. We observe that:
  • On the in-distribution GSM8K, direct distillation cuts tokens by over 50% (425 vs. 917) with only a minor accuracy drop (90.7% vs. 91.7%), showing effectiveness on simpler tasks.
  • On OOD tasks like MATH500, AIME24, and AMC, accuracy drops sharply despite modest token savings (e.g., 88.6% vs. 93.0% on MATH500), indicating poor generalization.
These results show that shorter CoTs alone do not generalize well. Direct distillation often omits critical steps, harming transfer. In contrast, DRP removes redundancy while preserving a step-wise, skill-aligned structure, enabling more robust reasoning. This suggests a key insight: preserving structural depth—even with moderate length—outperforms naive compression. See Appendix F for examples.